OpenStudy (anonymous):

How do I find the prime factorization for 180? I've looked in sundry places for how to find prime factorization, but alas, I continue to follow a confused path.

OpenStudy (unklerhaukus):

factor out one prime factor at a time if you have to 180=2*90=

hero (hero):

You just simply keep reducing the number in accordance with its factors: 180 = 2(90) = 2(2)(45) = 2(2)(3)(15) = 2(2)(3)(3)(5)